A rescue helicopter lifts a 90.4kg person straight up by the means of a cable. The person has an upward acceleration of 0.679 m/s2 and is lifted from rest through a distance of 12.2 m. a) What is the tension in the cable? How much work is done b) by the tension in the cable and c) The persons weight? d) Use the work-energy theorem And find the final speed of the person
Answer #1

A)

Tension, T = m(g + a)

T = 90.4 (9.8 + 0.679) = 947.30 N

B)

Work, W = Td

W = 947.3 x 12.2 = 11557.08 J

C)

Work, W = - 90.4 x 9.8 x 12.2 = - 10808.22 J

D)

Velocity, v = sqrt[2(11557.08 - 10808.22)/90.4] = 4.07 m/s

Comment in case any doubt please rate my answer....